A tow truck pulls a car from rest onto a level road. The tow truck exerts a horizontal force of 1500 N on the car. The frictional force on the car is 810 N. Calculate the work done by each of the following forces on the car as the car moves forward 12 m:
(a) the force of the tow truck on the car
(b) the force of friction
1
Expert's answer
2019-05-10T11:26:13-0400

The work done by definition

"W=Fd\cos\theta"


(a) The work done by the force of the tow truck on the car  

"W=1500\:\rm{N}\times 12\:\rm{m}\times \cos 0^{\circ}=18\: 000\:\rm{J}"

(b) The work done by the frictional force

"W=810\:\rm{N}\times 12\:\rm{m}\times \cos 180^{\circ}=-9\:720\:\rm{J}"